# Visual Semantic Navigation with Real Robots

### v1

@@ Abstract
robotics, embodied agents, vision-based navigation, artificial intelligence, reinforcement learning
[…]
[uah]organization=University of Alcala, Department of Signal Theory and Communications, city=Alcala de Henares, postcode=28801, country=Spain
@@ 1 Introduction
Is it possible for a robotic agent to navigate and interact in a real world environment the same way we humans do? This is the question that the embodied AI research community is pursuing to answer. The problem is formally known as Visual Semantic Navigation (VSN). The mimic of human navigation is a challenging task for robots, specially in unseen scenarios, since it requires an efficient exploration of the environment and an understanding of the objects and structures present in it. For unknown scenarios, humans can leverage prior semantic information achieved from previous scenes to navigate in new environments, but it is still a difficult task to incorporate that knowledge into embodied agents, especially in real robotic platforms. The possibilities of autonomous robots are endless, from assistive robots that can accompany a person with reduced mobility to a specific location, to platforms that can help in complex work scenarios, like search and rescue or logistic centers. Although there are several VSN works that achieve near-human navigation performance, e.g. (Ramrakhya et al., 2023), these systems are trained and tested in virtual environments, so how these models would behave in the real world is still a question to answer.
@@
In this work, we focus on embedding VSN models in real robotic platforms. We address the Object-Goal Navigation (ObjectNav(Batra et al., 2020) problem. As it is seen in figure 1, in ObjectNav task the agent has to navigate from a random spawn point to certain object goals present in the scene, mainly using vision-based sensors. In contrast with geometric navigation approaches, where the navigation problem is solved typically by using a map or generating it on the fly, VSN models are learning-based approaches that use no metric map. Therefore, VSN solutions must learn visual representations of the environment to reduce the exploration time and better generalize to unseen scenes and object categories. It is interesting to note that most of these solutions combine reinforcement learning (RL) strategies with recent advancements in deep learning models for visual perception so as to tackle the navigation problem in virtual environments where the embodied AI agents are trained and tested.
@@
    We first provide to the embodied agents community the ROS4VSN development, a novel Robot Operating System (ROS) (Quigley et al., 2009) based architecture that allows to test and compare different VSN models in real robots. Our ROS4VSN development is model agnostic, thus, any VSN solution can be integrated into it with ease. Section 3.1 includes all the technical details.
@@
    We also have embedded two state-of-the-art VSN models in two different robotic platforms. The selected models have been PIRLNav (Ramrakhya et al., 2023) and VLV (Chang et al., 2020). To achieve this integration, it is necessary to make technical adaptations to the models so that they transition from interacting with observations provided by simulation environments to observations from the real world. We detail all these technical modifications to the models in Section 3.2.
@@
    Finally, we provide a detailed experimental evaluation of the adapted VSN models, using our ROS4VSN, with two different robots, in a real world scenario (Section 4). The main question we want to address with the designed experiments is: Are the state-of-the-art VSN models able to successfully operate with real robots? We have measured their success rate in real world navigation experiments, which has allowed us to analyze the difference in performance compared to those tests in simulation environments, where embodied AI agents are typically tested. Our work shows that these VSN solutions perform noticeably differently when evaluated in real-world and simulated situations. With the ultimate goal of improving the performance and efficiency of VSN systems in real robots, we hope that our work will help to provide the groundwork for tackling this significant challenge.
@@ 2 Related Work
Visual Semantic Navigation. To navigate in unfamiliar environments, traditional methods use depth sensors (Newcombe et al., 2011; Thrun et al., 2001) and RGB cameras (Jones and Soatto, 2011; Sattler et al., 2018) to build geometric maps and simultaneously determine the robot’s position in relation to the map. This is known as Simultaneous Localization and Mapping (SLAM) (Abaspur Kazerouni et al., 2022; Campos et al., 2021; Labbé and Michaud, 2022)). Typically, these SLAM models use heuristic algorithms to create graph-based representations of the environment, allowing the robot to visit the different nodes of the graph when navigating to specific points. Semantic SLAM (e.g. (Zhang et al., 2018; Rosinol et al., 2020; Jin et al., 2023)) expands upon SLAM by integrating semantic data from the environment, allowing the robot to identify and store objects in memory.
@@
A recent approach, made possible by advances in machine learning and computer vision, involves designing navigation policies that directly train deep neural networks to learn semantic information from visual observations in an end-to-end fashion (e.g. (Ramrakhya et al., 2022; Yadav et al., 2023b; Gutiérrez-Maestro et al., 2019; Khandelwal et al., 2022; Chaplot et al., 2020; Chang et al., 2020)). This approach is termed visual semantic navigation (VSN). These models often rely on the use of CNNs as visual encoders followed by RNNs; that are in charge of predicting an action distribution directly from raw input observations. The neural networks are trained using imitation learning (IL) or reinforcement learning (RL) approaches.
@@
When IL is applied to the visual navigation problem, navigation policies are learnt from expert demonstrations (e.g. (Ramrakhya et al., 2022; Yadav et al., 2023b)). It can also be used combined with a RL finetuning phase to achieve better performance (Ramrakhya et al., 2023).
@@
Other works focus on the use of an end-to-end RL approach to solve ObjectNav navigation (Zhu et al., 2017; Gutiérrez-Maestro et al., 2019; Wijmans et al., 2020; Khandelwal et al., 2022; Liu et al., 2022; Yadav et al., 2023a). Some authors have proposed to combine the RL training with different strategies, like auxiliary tasks (Ye et al., 2021), improved visual representations via object relation graphs (Yang et al., 2018), semantic segmentations (Mousavian et al., 2018) or combining audio feedback with the visual inputs (Wang et al., 2023; Kondoh and Kanezaki, 2023).
@@
Modular-learning based approaches (Chaplot et al., 2020; Chang et al., 2020) decompose the navigation process in separate modules that execute different tasks. It is common for these methods to be composed of a high-level semantic exploration module trained by RL that indicates the agent sub-goals that have to be reached by a low-level navigation policy. Modular learning can be also combined with offline RL (Shah et al., 2022) techniques to leverage navigation behaviors from fixed datasets, without any additional online data collection or fine-tuning.
@@
Finally, there are different approaches that try to tackle the problem of rapidly adapting to unseen environments in visual navigation via meta-learning (Wortsman et al., 2019; Luo et al., 2021; Zhang et al., 2022). These methods are trained on a variety of different environments (usually designated as tasks) and are able to generalize to unseen environments by learning a policy that can be quickly adapted to new environments. And the recent progress in large language models (LLMs) has led to the possibility of using them to solve the visual navigation problem (Huang et al., 2023; Zhou et al., 2023) as well. In this case, the LLMs are used as a reasoning module in charge of understanding the semantic information present on the environment. They then share this information with different modules in charge of navigating to the specified goal.
@@
Our goal in this work consists in the technical embedding of different state-of-the-art VSN models in several real robots. We also provide the novel ROS4VSN package that allows to easily test and compare different VSN methods in the real world. To the best of our knowledge, our work is the first to develop a model agnostic ROS package for visual semantic navigation.
@@
Simulation-to-reality transfer. Deploying a model trained in simulation to a real robot is a challenging task. Due to logistical limitations, it is often not possible to train a model in the real world, especially when using RL techniques. For example Kim et al. (2022), propose a monocular vision-based time-to-collision estimation for small drones by domain adaptation of simulated images. Their method converts simulated images into real-like synthetic images using sim-to-real method. This is done with the aim of minimizing efforts and time invested in the collection of training datasets within real-world scenarios, while simultaneously maximizing the advantages inherent in simulated environments. Overall, it is necessary to develop methods that allow to efficiently transfer the knowledge learnt in simulation to the real world (Kadian et al., 2020). Different approaches have been proposed to solve this problem. For instance, CAD2RL (Sadeghi and Levine, 2017) system achieved remarkable success in training a collision avoidance policy entirely within a simulated environment. This breakthrough was subsequently tested on real aerial drones, with promising results.
@@
By focusing on simulation refinement (Son et al., 2020), the accuracy of simulations can be improved by exploiting the disparities between simulated and real-world observations. In the field of locomotion, training legged robotic systems in a simulated environment and subsequently transferring the acquired policies to real-world applications (Hwangbo et al., 2019; Agarwal et al., 2022) has always been a challenging task.
@@
Finally, the excellent study by Gervet et al. (2022) shows how different visual navigation methods perform in the real-world. Indeed, this work is closely related to our research. However, we would like to highlight the novel contributions that our work offers. While (Gervet et al., 2022) focuses mainly on the comparison of navigation methods, we here, along with a similar study, release to the robotics community the modular ROS4VSN architecture. Our main goal is to facilitate the prototyping of new VSN solutions on real robots. So, we offer a ROS compatible software architecture, model agnostic, that allows a simple integration of different VSN approaches in ROS robots. Furthermore, we include in our study more recent VSN solutions than the ones reported in (Gervet et al., 2022), as the PIRLNav model (Ramrakhya et al., 2023), which defines the state-of-the-art for the ObjectNav problem. We also provide, for the first time, a detailed analysis on how a model directly trained with real videos, such as the VLV (Chang et al., 2020), performs in real robots. This allows us to compare, as in (Gervet et al., 2022), how a modular-learning model (i.e. VLV (Chang et al., 2020)), compares with a typical end-to-end learning approach (i.e. PIRLNav). Interestingly, our study also concludes, like in (Gervet et al., 2022), that modular-learning approaches perform better in the real world. Our results, with two different robots confirm that. Overall, we hope that our ROS library will help to further advance the field of visual semantic navigation in real robots.
@@ 3 Methodology
In this research work, our main objective has been to efficiently integrate various state-of-the-art VSN models on multiple robotic platforms. To fulfill our objectives, the first step of the proposed methodology has been to develop a ROS-based solution to ease the integration of VSN solutions into real robots. It is crucial that the approach is model-agnostic, allowing the integration of any VSN-designed model. We have named this development ROS4VSN, and it is detailed in Section 3.1. Once the ROS4VSN system is available, it is time to select state-of-the-art VSN models that will allow us to experimentally evaluate them with real robots. In Section 3.2, we justify the selected VSN models and detail the technical modifications made to integrate them into ROS4VSN.
@@ 3.1 ROS4VSN: ROS for Visual Semantic Navigation
We have designed ROS4VSN library to be modular and flexible, so that it can be easily adapted to different robots and VSN models. It is built on top of ROS Noetic (Quigley et al., 2009) open-source robotic middleware, because its flexibility, support, compatibility and popularity among the robotics community. ROS provides a collection of useful tools, libraries, and conventions to simplify the task of creating complex and robust robot behaviors across a wide variety of robotic platforms, which makes it perfect for our framework.
We design the architecture of the framework, so it has three main capabilities: it can receive and process information from the environment, infer actions using an AI VSN model, and control the actuators of a platform to reach a specific navigation goal. It makes it easy to integrate different VSN models, since it only needs to replace the model with which the experiments are to be carried out.
The architecture is divided into the following main packages, each of which plays a specific role: robot_api, camera_api, discrete_move and visual_semantic_navigation. These packages are connected to each other through ROS topics and services, and also to external hardware devices: a camera (RGB + Depth) and a differential drive robotic platform.
Figure 2 shows a visual representation of the global architecture scheme, illustrating the connections between the different developed ROS packages and the hardware devices.
Figure 2: Architecture scheme of the ROS44VSN framework. It shows the different packages, their topics and the connections within them and the hardware devices.
@@ 3.1.1 Robot API
This package is responsible for controlling the actuators of the robot and sending odometry information to the discrete_move and visual_semantic_navigation packages (depicted in figure 2). It is typically designed by the manufacturer of the robot, so it can be different depending on the employed platform. In our particular case, the development of the framework and experiments were done using two robots (see Figure 8). First, a Turtlebot 2 robot, so the standard turtlebot2 (Ltd., 2023a) package is integrated as robot_api. Since the Turtlebot 2 expects the velocity commands in the /mobile_base/commands/velocity topic, we perform a remap to the /cmd_vel topic from our discrete_move package. Our second robot is know as LOLA2 (Nasri et al., 2022), for which we have developed the complete robot_api package, so as to guarantee the compatibility with the rest of ROS4VSN architecture.
@@ 3.1.2 Camera API
This package is responsible for capturing RGB and depth images from the camera. It publishes them through the /camera/color and /camera/depth topics, respectively. The camera used in our robots is the Orbecc Astra S, a RGB plus depth camera, based on structured light technology. We have adapted the official ros_astra_camera package (Ltd., 2023b) to be integrated in our ROS4VSN architecture. The modular design of ROS4VSN allows it to be used with any other type of camera on the robots, simply by adapting this Camera API package.
@@ 3.1.3 Discrete Move package
This package has been developed with the purpose of providing a precise and customizable control of the robotic platform through discrete navigation commands. Note that this is the way most VSN models interact with their agents (e.g. (Ramrakhya et al., 2023; Chang et al., 2020)). In embodied AI, navigation in simulated environments is performed using discrete action commands that agents execute to reach the specific goals: move forward (25 cm), turn left or right (30 degrees), or stop. Our ROS4VSN package acts as a server to which clients can request a set of discrete movements and configure the forward distance or turning angles. The package communications scheme is shown in detail figure 3. It is in charge of executing the actions requested by the visual_semantic_navigation package and sending a response when the action is completed.
@@
The set of movements allowed by the package consists of the following actions: turn_left, turn_right, move_forward, move_backward and stop. All the actions are fully customizable in terms of distance and angle, except for the stop action, which does not require any additional parameters since it just stops the robot. This package has been designed as a ROS service, so the communication between the visual_semantic_navigation package and the discrete_move package is done synchronously and bidirectionally. That way, the visual_semantic_navigation package can wait for the response of the discrete_move package when the action has been completed before sending any new action request. The discrete_move server is in charge of sending the right /cmd_vel commands to the robot_api package, so the robot can execute the requested action and to receive the /odometry topic information from the robot_api package, so it can calculate the movement done by the robot, and stop the action when the requested action has been finished.
Embodied AI navigation environments, such as Habitat (Szot et al., 2021), are simulation environments where there are no movement errors in the agents. However, our scenario is the real world, with real robots, therefore ROS4VSN must integrate error control strategies. To achieve this, the discrete_move package includes two error correction strategies: one for the turn error $\epsilon_{turn}$ and one for the move straight error $\epsilon_{straight}$.
@@
The smoothness of the movement and the time needed to complete it depend on the percentage of the path in which the robot is accelerating and decelerating, as well as on the initial speed. By properly adjusting these parameters, a smoother movement and a more efficient navigation time can be achieved. Figure 4 illustrates the total distance that the robot must travel for a move_forward (or move_backward) command. This distance is divided so that the robot performs the following phases: acceleration, displacement at constant speed and deceleration. The instantaneous speed of the robot is controlled so that during the first phase, there is a uniformly accelerated motion, according to the following equation: $v=\sqrt{v_{init}^{2}+2a\epsilon_{straight}}$, where $a$ is the desired acceleration, $v_{init}$ represents the initial speed of the robot, and $\epsilon_{straight}$ encodes the distance covered by the robot. Note that we must continuously read $\epsilon_{straight}$ using our ROBOT API, in order to dynamically adjust the speed. For the deceleration stage, we employ an equivalent negative acceleration in the previous equation. With these equations we can progressively adjust the linear speed ($v$) that is sent to the platform in order to obtain an smooth navigation.

### v2

@@ Abstract
robotics; embodied agents; vision-based navigation; artificial intelligence; reinforcement learning
@@ 1 Introduction
Can a robotic agent navigate and interact in the real world as seamlessly as humans do? This is the fundamental question driving research within the embodied AI community. The problem is formally known as Visual Semantic Navigation (VSN), e.g [1, 2, 3]. However, mimicking human navigation is a challenging task for robots, particularly in unseen environments, as it requires efficient exploration and a deep understanding of the objects and structures within the space. For unknown scenarios, humans can leverage prior semantic information achieved from previous scenes to navigate in new environments, but it is still a challenging task to incorporate that knowledge into embodied agents, especially in real robotic platforms navigating in the real world. The potential of autonomous robots with these advanced navigation capabilities is vast, ranging from assistive robots that can guide individuals with reduced mobility to specific locations, to platforms that can aid in complex environments such as search and rescue operations or logistic centers.
Technically, in this work, we focus on embedding VSN models in real robotic platforms. This is our main objective. We focus on the Object-Goal Navigation (ObjectNav[4] problem. As it is seen in figure 1, in ObjectNav task the agent has to navigate from a random position to certain object goals present in the scene, mainly using vision-based sensors. In contrast with traditional geometric navigation approaches, where the navigation problem is solved typically by using a map or generating it on the fly (e.g. SLAM), VSN models are learning-based approaches that use no metric map. Therefore, VSN solutions must learn visual representations of the environment to reduce the exploration time and better generalize to unseen scenes and object categories. Notably, many of these solutions combine reinforcement learning (RL) and/or imitation learning (IL) strategies with recent advances in deep learning models for visual perception to address the navigation problem in virtual environments, where embodied AI agents are trained and tested. However, we believe it is crucial to thoroughly investigate how the latest solutions to the VSN problem perform when deployed with real robots navigating in real-world environments. This is where our work makes its most significant contribution. Although there are VSN works that achieve near-human navigation performance, e.g [1], these systems are mainly trained and tested in virtual environments, so how these models would behave in the real world is still a question to answer.
@@
    We first provide to the embodied agents community the ROS4VSN development, a novel Robot Operating System (ROS) [5] based architecture that allows testing and comparing different VSN models in real robots. Our ROS4VSN development is model agnostic; thus, any VSN solution can be integrated into it with ease. Section 3.1 includes all the technical details.
@@
    We also have embedded two state-of-the-art VSN models in two different robotic platforms. The selected models have been PIRLNav [1] and VLV [3]. To achieve this integration, it is necessary to make technical adaptations to the models so that they transition from interacting with observations provided by simulation environments to observations from the real world. We detail all these technical modifications to the models in Section 3.2.
@@
    Finally, we propose an experimental evaluation of the adapted VSN models, using our ROS4VSN, with two different robots, in a real world scenario (Section 4). The main question we want to address with the designed experiments is: Are the state-of-the-art VSN models able to successfully operate with real robots? We have measured their success rate in real world navigation experiments, which has allowed us to analyze the difference in performance compared to those tests in simulation environments, where embodied AI agents are typically tested. Our work shows that these VSN solutions perform noticeably differently when evaluated in real-world and simulated situations. With the ultimate goal of improving the performance and efficiency of VSN systems in real robots, we hope that our work will help to provide the groundwork for tackling this significant challenge.
@@ 2 Related Work
Visual Semantic Navigation. To navigate in unfamiliar environments, traditional methods use depth sensors [6, 7] and RGB cameras [8, 9] to build geometric maps and simultaneously determine the robot’s position in relation to the map. This is known as Simultaneous Localization and Mapping (SLAM) [10, 11, 12]). Typically, these SLAM models use heuristic algorithms to create graph-based representations of the environment, allowing the robot to visit the different nodes of the graph when navigating to specific points. Semantic SLAM (e.g [13, 14, 15]) expands upon SLAM by integrating semantic data from the environment, allowing the robot to identify and store objects in memory.
@@
A recent approach, made possible by advances in machine learning and computer vision, involves designing navigation policies that directly train deep neural networks to learn semantic information from visual observations in an end-to-end fashion (e.g [16, 17, 18, 19, 20, 3]). This approach is termed visual semantic navigation (VSN). These models often rely on the use of CNNs as visual encoders followed by RNNs; that are in charge of predicting an action distribution directly from raw input observations. The neural networks are trained using imitation learning (IL) or reinforcement learning (RL) approaches.
@@
When IL is applied to the visual navigation problem, navigation policies are learnt from expert demonstrations (e.g [16, 17]). It can also be used combined with an RL fine-tuning phase to achieve better performance [1].
@@
Other works focus on the use of an end-to-end RL approach to solve ObjectNav navigation [21, 18, 22, 19, 23, 24]  [25, 26] . Some authors have proposed combining the RL training with different strategies, like auxiliary tasks [27], improved visual representations via object relation graphs [28], semantic segmentations [29] or combining audio feedback with the visual inputs [30, 31].
@@
Modular-learning based approaches [20, 3, 32, 33, 34]  [2, 35, 36, 37, 38] decompose the navigation process in separate modules that execute different tasks. It is common for these methods to be composed of a high-level semantic exploration module trained by RL that indicates the agent subgoals that have to be reached by a low-level navigation policy. Modular learning can be also combined with offline RL [39] techniques to leverage navigation behaviors from fixed datasets, without any additional online data collection or fine-tuning.
@@
Finally, there are different approaches that try to tackle the problem of rapidly adapting to unseen environments in visual navigation via meta-learning [40, 41, 42]. These methods are trained on a variety of different environments (usually designated as tasks) and are able to generalize to unseen environments by learning a policy that can be quickly adapted to new environments. And the recent progress in large language models (LLMs) has led to the possibility of using them to solve the visual navigation problem [43, 44] as well. In this case, the LLMs are used as a reasoning module in charge of understanding the semantic information present on the environment. They then share this information with different modules in charge of navigating to the specified goal.
@@
Our goal in this work is not to develope VSN approaches, but to integrate various state-of-the-art VSN models into multiple real-world robots by using our novel ROS4VSN library. Technically, we have chosen to integrate the PIRLNav [1] and VLV [3] models into two different robots. These integrations required several technical adaptations, particularly in the areas of sensor data integration and navigation planning. Overall, we are able to show how ROS4VSN allows easily testing and comparing different VSN methods in the real world. To the best of our knowledge, our work is the first to develop a model agnostic ROS package for visual semantic navigation, where multiple models can be integrated.
@@
Simulation-to-reality transfer in robotic navigation. Deploying a model trained in simulation to a real robot is a challenging task. Due to logistical constraints, training a model in the real world —especially with RL techniques— is often impractical, prompting the use of alternative methods to address this challenge For example, Kim et al. [45] propose a monocular vision-based time-to-collision estimation for small drones by domain adaptation of simulated images. Their method converts simulated images into real-like synthetic images using a sim-to-real method. This is done with the aim of minimizing efforts and time invested in the collection of training datasets within real-world scenarios, while simultaneously maximizing the advantages inherent in simulated environments.
@@
Overall, it is necessary to develop methods that allow to efficiently transfer the knowledge learnt in simulation to the real world [46]. Different approaches have been proposed to solve this problem. For instance, CAD2RL [47] system achieved remarkable success in training a collision avoidance policy entirely within a simulated environment. This breakthrough was subsequently tested on real aerial drones, with promising results. By focusing on simulation refinement [48], the accuracy of simulations can be improved by exploiting the disparities between simulated and real-world observations. In the field of locomotion, training legged robotic systems in a simulated environment and subsequently transferring the acquired policies to real-world applications [49, 50] has always been a challenging task.
@@
For the problem of VSN, we have the study by Gervet et al. [51] that shows how their approaches perform in real-world settings. However, we would like to highlight the novel contributions that our work offers. First, while [51] focuses mainly on the comparison of their navigation methods, we here, along with a similar study, release to the research community the modular ROS4VSN software architecture. Our main goal is to facilitate the prototyping of new VSN solutions on real robots. So, we offer a ROS-compatible software architecture, model agnostic, that allows a simple integration of different VSN approaches in ROS robots. In this way, future VSN solutions will be able to be tested on real robots in a convenient and straightforward manner. Second, we include in our study more recent VSN solutions than the ones reported in [51], as the PIRLNav model [1], which defines the state-of-the-art for the ObjectNav problem. Third, we also provide, for the first time, a detailed analysis on how a model directly trained with real videos, such as the VLV [3], performs in real robots. This allows us to compare, as in [51], how a modular-learning model (i.e VLV [3]), compares with a typical end-to-end learning approach (i.e PIRLNav). Interestingly, our study also concludes, like in [51], that modular-learning approaches perform better in the real world. Fourth, in our study, we employed two different robotic platforms: one commercially available and widely used by various laboratories, and another custom-built. This demonstrates the versatility of the proposed solution, showing that it can be integrated into different robots. And finally, in our work, we propose an experimental evaluation specifically designed for testing in the real world, which can be employed in future research studies. Overall, we hope that our ROS-based library will help to further advance the field of visual semantic navigation in real robots.
@@ 3 Methodology
In this research work, our main objective has been to efficiently integrate various state-of-the-art VSN models on multiple robotic platforms. To fulfill our goals, the first step of the proposed methodology has been to develop a ROS-based solution to ease the integration of VSN solutions into real robots. It is crucial that the approach is model-agnostic, allowing the integration of any VSN-designed model. We have named this development ROS4VSN, and it is detailed in Section 3.1. Once the ROS4VSN system is available, we need to select some state-of-the-art VSN models that will allow us to experimentally evaluate them with real robots. In Section 3.2, we justify the selected VSN models and detail the technical modifications made to integrate them into ROS4VSN.
@@ 3.1 ROS4VSN: ROS for Visual Semantic Navigation
We have designed ROS4VSN library to be modular and flexible, so that it can be easily adapted to different robots and VSN models. It is built on top of ROS Noetic [5] open-source robotic middleware, because of its flexibility, support, compatibility and popularity among the robotics community. ROS provides a collection of useful tools, libraries, and conventions to simplify the task of creating complex and robust robot behaviors across a wide variety of robotic platforms, which makes it perfect for our framework. We design the architecture of the framework, so it has three main capabilities: it can receive and process information from the environment, infer actions using an AI VSN model, and control the actuators of a platform to reach a specific navigation goal. It makes it easy to integrate different VSN models, since it only needs to replace the model with which the experiments are to be carried out. The architecture is divided into the following main packages, each of which plays a specific role: robot_api, camera_api, discrete_move and visual_semantic_navigation. These packages are connected to each other through ROS topics and services, and also to external hardware devices: a camera (RGB + Depth) and a differential drive robotic platform. Figure 2 shows a visual representation of the global architecture scheme, illustrating the connections between the different developed ROS packages and the hardware devices.
Figure 2: Architecture scheme of the ROS44VSN framework. It shows the different packages, topics and connections within them and the hardware devices.
@@ 3.1.1 Robot API
This package is responsible for controlling the actuators of the robot and sending odometry information to the discrete_move and visual_semantic_navigation packages (depicted in figure 2). It is typically designed by the manufacturer of the robot, so it can be different depending on the employed platform. In our particular case, the development of the framework and experiments were done using two robots (see Figure 8). First, a Turtlebot 2 robot, so the standard turtlebot2 [52] package is integrated as robot_api. Since the Turtlebot 2 expects the velocity commands in the /mobile_base/commands/velocity topic, we perform a remapping to the /cmd_vel topic from our discrete_move package. Our second robot is known as LOLA2 [53]. We have developed its complete robot_api package to guarantee the compatibility with the rest of ROS4VSN architecture.
@@ 3.1.2 Camera API
This package is responsible for capturing RGB and depth images from the camera. It publishes them through the /camera/color and /camera/depth topics, respectively. The camera used in our robots is the Orbecc Astra S, an RGB plus depth camera, based on structured light technology. We have adapted the official ros_astra_camera package [54] to be integrated in our ROS4VSN architecture. The modular design of ROS4VSN allows it to be used with any other type of camera on the robots, simply by adapting this Camera API package.
@@ 3.1.3 Discrete Move package
This package has been developed with the purpose of providing a precise and customizable control of the robotic platform through discrete navigation commands. Note that this is the way most VSN models interact with their agents (e.g [1, 3]). In embodied AI, navigation in simulated environments is performed using discrete action commands that agents execute to reach the specific goals: move forward (25 cm), turn left or right (30 degrees), or stop. Our ROS4VSN package acts as a server to which clients can request a set of discrete movements and configure the forward distance or turning angles. The package communications scheme is shown in detail figure 3. It is in charge of executing the actions requested by the visual_semantic_navigation package and sending a response when the action is completed.
@@
The set of movements allowed by the package consists of the following actions: turn_left, turn_right, move_forward, move_backward and stop. All the actions are fully customizable in terms of distance and angle, except for the stop action, which does not require any additional parameters since it just stops the robot. This package has been designed as a ROS service, so the communication between the visual_semantic_navigation package and the discrete_move package is done synchronously and bidirectionally. That way, the visual_semantic_navigation package can wait for the response of the discrete_move package when the action has been completed before sending any new action request. The discrete_move server is in charge of sending the right /cmd_vel commands to the robot_api package, so the robot can execute the requested action and receive the /odometry topic information from the robot_api package. That way, it can calculate the movement done by the robot, and stop the action when the requested action has been finished.
Embodied AI navigation environments, such as Habitat [55], are simulation environments where there are no movement errors in the agents. However, our scenario is the real world, with real robots. Therefore, ROS4VSN must integrate error control strategies. To achieve this, the discrete_move package includes two error correction strategies: one for the turn error $\epsilon_{turn}$ and one for the move straight error $\epsilon_{straight}$.
@@
The smoothness of the movement and the time needed to complete it depend on the percentage of the path in which the robot is accelerating and decelerating, as well as on the initial speed. By properly adjusting these parameters, a smoother movement and a more efficient navigation time can be achieved. Figure 4 illustrates the total distance that the robot must travel for a move_forward (or move_backward) command. This distance is divided so that the robot performs the following phases: acceleration, displacement at constant speed and deceleration. The instantaneous speed of the robot is controlled so that during the first phase, there is a uniformly accelerated motion, according to the following equation: $v=\sqrt{v_{init}^{2}+2a\epsilon_{straight}}$, where $a$ is the desired acceleration, $v_{init}$ represents the initial speed of the robot, and $\epsilon_{straight}$ encodes the distance covered by the robot. Note that we must continuously read $\epsilon_{straight}$ using our ROBOT API, in order to dynamically adjust the speed. For the deceleration stage, we employ an equivalent negative acceleration in the previous equation. With these equations, we can progressively adjust the linear speed ($v$) that is sent to the platform to obtain a smooth navigation.